# Codec Does Matter: Exploring the Semantic Shortcoming of Codec for Audio Language Model

### v2

@@ Abstract
Introduction
In recent years, Large Language Models (LLMs) such as GPT (Brown et al. 2020) have demonstrated remarkable capabilities in modeling complex, high-dimensional data across various domains, including text and image generation (Zhao et al. 2023; Liu et al. 2024a). Inspired by these successes, there has been significant interest (Agostinelli et al. 2023; Borsos et al. 2023; Wang et al. 2023; Yang et al. 2023b) in exploring the application of LLMs to audio generation.
Audio codecs (Zeghidour et al. 2021) have emerged as a critical technique for audio LLMs, bridging the gap between continuous audio waveforms and token-based language models. By discretizing high-rate audio signals into a finite set of tokens, these codecs enable the application of LLM architectures to audio data, leveraging the successes of textual LLMs.
However, prior research on audio codecs has primarily focused on achieving lower compression rates and higher reconstruction quality (Kumar et al. 2024; Défossez et al. 2022; Yang et al. 2023a). Meanwhile, many efforts in audio generation have concentrated on enhancing model architecture, scaling, or leveraging larger datasets. For instance, AudioLM (Borsos et al. 2023) adopts a two-stage pipeline that models the acoustic token in an autoregressive way conditioned on the semantic token. VALL-E (Wang et al. 2023), the first TTS framework to leverage large, diverse, and multi-speaker speech data, demonstrates strong in-context learning capabilities similar to GPT-3, treating TTS as a language modeling task on audio codecs. MusicGen (Copet et al. 2024) generates music using a single-stage transformer LM alongside efficient token interleaving patterns. Similarly, UniAudio (Yang et al. 2023b) scaled up to 165K hours of audio and 1B parameters, utilizing LLM techniques to generate tokens for various types of audio, including speech, sounds, music, and singing, given different input conditions.
While these works have shown success in developing audio language models, they all rely on the acoustic codecs such as Encodec (Défossez et al. 2022) or Soundstream (Zeghidour et al. 2021) for audio tokenization and de-tokenization. However, these acoustic codecs were originally designed for audio compression rather than for audio language models. This misalignment means the design may not be optimal for audio language modeling.
@@ Introduction
To address this issue, approaches like SpeechTokenizer (Zhang et al. 2023) have attempted to disentangle speech into separate tokens for content and timbre and perform distillation-based semantic and acoustic integration. However, this method may not integrate smoothly with all audio LLMs, especially those requiring uniform token treatment across different layers, such as utilizing flattened codec tokens (Yang et al. 2023b; Copet et al. 2024).
@@
Related Works
Audio Language Model
The success of Large Language Models (LLMs) has sparked a significant trend in leveraging language foundation models for audio generation tasks (Rubenstein et al. 2023; Zhang et al. 2024; Wu et al. 2023b, a; Yang et al. 2023b; Chen et al. 2023). Audio, much like language, consists of variable-length sequences, making it well-suited for modeling with language foundation models. One pioneering method, AudioLM (Borsos et al. 2023), employs a multi-stage strategy to harness the predictive capabilities of foundation models for generating tokens unconditionally. This approach involves predicting semantic tokens from various conditions (e.g., phonemes, text descriptions, MIDI) in the initial stage, followed by transforming them into acoustic tokens through coarse-to-fine modeling, ultimately generating the waveform. Representative systems such as SPEAR-TTS (Kharitonov et al. 2023) for speech synthesis and MusicLM (Agostinelli et al. 2023) for music generation have also been proposed. However, the two-stage process can lead to complexity in training and suboptimal performance due to the separate development of semantic and acoustic tokens, leading to error accumulation.
Conversely, recent advancements have shown that methods employing a single-stage language model outperform two-stage approaches. For example, VALL-E (Wang et al. 2023) utilizes an autoregressive (AR) model to predict the first token and a non-autoregressive (NAR) model to estimate the residual tokens, demonstrating superior performance compared to AudioLM. Similarly, MusicGen (Copet et al. 2024) employs a single-stage transformer language model and incorporates a delay pattern strategy for efficient token interleaving, achieving better results than MusicLM. Other notable works include CLAM-TTS (Kim et al. 2024), VoiceCraft (Peng et al. 2024), and UniAudio (Yang et al. 2023b).
@@
Audio Codec
Recent advancements have seen a surge in deep learning methodologies employing vector quantization (Van Den Oord, Vinyals et al. 2017) to reconstruct continuous signals into discrete representations for AR generation. Notably, audio codecs based on the VQ-GAN framework (Esser, Rombach, and Ommer 2021) have gained prominence. For example, SoundStream (Zeghidour et al. 2021) introduces a versatile codec adaptable to various audio types, integrating Residual Vector Quantization (RVQ) and Generative Adversarial Network (GAN) to refine quantization and reconstruction. Similarly, Encodec (Défossez et al. 2022) enhances compression through a multi-scale discriminator and a loss-balancing strategy alongside a language model. HiFi-Codec (Yang et al. 2023a) employs Group-Residual Vector Quantization (GRVQ) to minimize the need for extensive codebooks while maintaining high reconstruction fidelity. DAC (Kumar et al. 2024) addresses codebook collapse, where some codes remain unused, by applying improved codebook learning to achieve higher compression rates.
These codecs primarily focus on acoustic reconstruction and higher compression rates, often overlooking their potential as tokenizers for audio LLMs. Some attempts have been made to develop more suitable tokenizers for audio LLMs. For example, SpeechTokenizer (Zhang et al. 2023) utilizes HuBERT to separate speech into distinct VQ components for content and timbre/acoustic details. This separation improves the modeling of content in the AR stage of VALL-E, while the NAR stage enriches the acoustic details. However, a distillation framework is exploited, this makes SpeechTokenizer may not be compatible with all LLM architectures, especially those that require uniform treatment of tokens, such as methods using flattened codec tokens (Yang et al. 2023b; Copet et al. 2024). Another attempt is presented by SemantiCodec (Liu et al. 2024b), which employs a pre-trained AudioMAE (Huang et al. 2022) to generate distinct semantic and acoustic tokens from mel-spectrograms. However, this method inherits the issues of SpeechTokenizer and introduces additional complexity in token modeling. Moreover, since the audioMAE is performed on 2D time-frequency mel-spectrograms, LLMs must effectively handle dual scales (time and frequency), which may require significant modifications to existing LLM structures.
@@ Methods
Acoustic Audio codec
As illustrated in Figure 1, our model builds upon the framework established by existing acoustic codecs such as Encodec (Défossez et al. 2022) and DAC(Kumar et al. 2024). An acoustic audio codec is composed of three main components: an acoustic encoder, a quantizer, and an acoustic decoder. The input of the codec is the raw waveform $\textbf{X}\in\mathbb{R}^{n}$, where $n$ represents the number of waveform samples. This waveform is fed into the acoustic encoder, which consists of several convolutional layers and employs temporal downscaling to extract frame-level latent acoustic features $\textbf{A}\in\mathbb{R}^{H_{a}\times T}$, where $H_{a}$ denotes the hidden size of the acoustic features and $T$ is the number of frames. These continuous features are then transformed into a series of discrete tokens $\textbf{Q}\in\mathbb{R}^{M\times T}$ using a Residual Vector Quantizer (RVQ) with $M$ quantizer layers. During training, a specific codebook for the quantizer is learned, enabling the conversion of discrete tokens back to continuous features $\textbf{A}_{q}\in\mathbb{R}^{H_{a}\times T}$. The acoustic decoder then reconstructs the waveform $\hat{\textbf{X}}$ from $\textbf{A}_{q}$ using several convolutional layers and temporal upsampling. The training process is supervised using various losses, including mel loss, STFT loss, and GAN loss, to ensure high-quality acoustic reconstruction.
Analysing Semantic Shortcoming
@@
Designing Auxiliary Semantic Module
Our approach employs a straightforward method that enhances audio codecs by directly concatenating semantic and acoustic features. Initially, we extract the semantic feature vector $\textbf{S}^{*}\in\mathbb{R}^{H_{s}\times T}$ from the audio waveform x. This extraction utilizes a self-supervised, pre-trained model such as HuBERT (Hsu et al. 2021) or wav2vec 2.0 (Baevski et al. 2020). The extracted features are then processed through multiple convolutional layers within a semantic encoder to yield the refined semantic feature vector S. Concurrently, the acoustic branch produces the feature A. These outputs, S and A, are subsequently concatenated using a linear projection $\phi$, formulated as:
@@
  • EnCodec  [^2] https://huggingface.co/facebook/encodec“˙24khz: The open-source EnCodec model (Défossez et al. 2022), trained on a diverse range of 24kHz audio data, can compress audio to bitrates between 1.5 and 24.0 kbps while maintaining high fidelity.
  • DAC  [^3] https://github.com/descriptinc/descript-audio-codec: The open-source DAC model (Kumar et al. 2024) utilizes enhanced VQ techniques. For our experiments, we employ the official 16kHz version.
  • SpeechTokenizer  [^4] https://github.com/ZhangXInFD/SpeechTokenizer: This model (Zhang et al. 2023) is a unified speech tokenizer that leverages distinct VQ layers to separate speech into content and timbre components. We utilize their official checkpoints in our evaluations.
Training Details of X-Codec
@@
Training Details of VALL-E
@@
  • WER (Word Error Rate): We utilize an Automatic Speech Recognition (ASR) model to transcribe the generated audio (Wang et al. 2023). The discrepancies between these transcriptions and the original texts are quantified using WER, providing a critical measure of audio intelligibility.
  • Sim-O (Similarity Objective): This metric assesses the objective similarity between synthesized speech and the original reference speech. Sim-O uses feature embeddings extracted from a pre-trained speaker verification model to measure this similarity (Hsu et al. 2021; Kim et al. 2024) [^7] https://github.com/microsoft/UniSpeech/tree/main/downstreams/speaker˙verification, reflecting the codec’s ability to preserve speaker characteristics.
  • UTMOS: We evaluate the audio quality using UTMOS, a Speech MOS (Mean Opinion Score) predictor (Saeki et al. 2022) [^8] https://github.com/tarepan/SpeechMOS that automatically measures the naturalness of speech. This metric provides insights into the overall auditory quality of the synthesized speech.
@@
Encodec (Défossez et al. 2022) C+DNS+AS+FSD+J 47.17 0.09 1.24 6.37 0.33 3.02
DAC (Kumar et al. 2024) C+DNS+V+AS+J+M 85.55 0.03 1.24 6.81 0.34 3.31
Speechtokenizer (Zhang et al. 2023) LibriSpeech 7.53 0.10 1.26 5.24 0.36 3.84
@@
Zero-shot TTS Results
We use librispeech-test-clean (Panayotov et al. 2015)for zero-shot TTS evaluation following VALL-E-continual-setting (Wang et al. 2023). The results in Table 1 demonstrate the following key findings:
@@
hubert-ls-960 (Hsu et al. 2021) - 3.3 4.1
Encodec (Défossez et al. 2022) 1 21.5 28.3
Encodec (Défossez et al. 2022) 8 17.5 27.0
DAC (Kumar et al. 2024) 1 26.3 32.7
DAC (Kumar et al. 2024) 12 21.7 33.2
Speechtoknizer (Zhang et al. 2023) 1 3.5 4.3
Speechtoknizer(Zhang et al. 2023) 8 3.6 4.5
@@
Analysing the Effect of Codec
To further analyse the above results caused by different audio codecs, we evaluate phonetic discriminability using the ABX error rate (Schatz et al. 2013). This metric assesses how well different codecs can distinguish between similar phonetic sounds within and across various contexts. We specifically examine the continuous representations for VQ as indicated by the results in the following table 2. We compare the performance of various models in terms of within and across phonetic discriminability:
@@
Music and Sound Generation
To the best of our knowledge, this is the first exploration into the potential benefits of incorporating semantic information into audio codecs for enhancing music and general sound generation through audio LLMs. Conventional methods for general audio representation learning, aiming at capturing the semantic discriminability of audios, are generally based on 2D mel-spectrogram, such as AudioMAE (Huang et al. 2022) and Beats (Chen et al. 2022). These methods are in stark contrast to traditional codecs that process audio sequentially, frame-by-frame. This difference poses challenges for direct integration into existing audio generation frameworks.
@@
Training Details of Self-Supervised General Audio Representation
HuBERT-General-Audio is trained using 8 NVIDIA H800 GPUs on 2.6 million tokens across 325,000 iterations. For training stability, we adopt an inverse square root learning schedule, a modification from the polynomial decay schedule originally utilized in (Hsu et al. 2021). The learning rate is set at 0.0003 with warmup steps of 32,000. Unlike the original HuBERT, which utilizes MFCCs as the training target unit designed specifically for speech, our model leverages the first VQ layer of Encodec as the training target for acoustic unit discovery in the general audio. This choice eliminates the need for the K-means discretization step, saving significant time and computational resources.
@@
Music Continuation
Training Details: Acquiring high-quality text-music pair data is challenging; therefore, we gathered approximately 100,000 hours of music-only data, including about one million songs for the music continuation task. We deployed nanoGPT  [^9] https://github.com/karpathy/nanoGPT to implement a GPT-2-medium (approximately 300M parameters) (Radford et al. 2019) as our generative model. This model utilizes the first VQ from our codec to construct the training sequences, with additional experiments involving multiple VQs detailed in the appendix. We set the block size of sequence modelling to 4096, corresponding to roughly 82 seconds of audio, and adjust the vocabulary size from 50,257 to 1024, matching our codec’s codebook size. Other training hyperparameters are consistent with previous GPT-2-medium configurations. We train 300,000 steps on 8 NVIDIA H800 GPUs. The batch size is set to 20, with a learning rate of 3e-4 and a warmup phase of 2000 steps.
Experiments: For music continuation, we randomly crop 600 samples with each 40 seconds in duration from the MUSDB18 dataset (Rafii et al. 2017). The initial 10 seconds of each sample are used as prompts for the audio LLM, while the subsequent 30 seconds are generated by the model. These generated segments are then compared against the corresponding ground truth (GT) segments. To ensure that the assessment is independent of the codec’s reconstruction fidelity, both the generated and GT audio are reconstructed using the first VQ layer of the codec, ensuring performance differences attributed solely to the generative models themselves.
The evaluation metrics of the generated music include: Frechet Distance (FD) computed using features from Pretrained Audio Neural Networks (PANNs) (Kong et al. 2020), Frechet Audio Distance (FAD), and FD-MERT Layer 9 (Li et al. 2023). The results, as summarized in Table 7, reveal that the X-Codec significantly outperforms the baseline acoustic codec across all metrics. This superior performance indicates the X-Codec has a better understanding and enabling more effective reproduction of complex musical structures.
Text-to-Sound

### v3

@@ Abstract
1 Introduction
In recent years, Large Language Models (LLMs) such as GPT [1] have demonstrated remarkable capabilities in modeling complex, high-dimensional data across various domains, including text and image generation [2, 3]. Inspired by these successes, there has been significant interest [4, 5, 6, 7] in exploring the application of LLMs to audio generation.
Audio codecs [8] have emerged as a critical technique for audio LLMs, bridging the gap between continuous audio waveforms and token-based language models. By discretizing high-rate audio signals into a finite set of tokens, these codecs enable the application of LLM architectures to audio data, leveraging the successes of textual LLMs.
However, prior research on audio codecs has primarily focused on achieving lower compression rates and higher reconstruction quality [9, 10, 11]. Meanwhile, many efforts in audio generation have concentrated on enhancing model architecture, scaling, or leveraging larger datasets. For instance, AudioLM [5] adopts a two-stage pipeline that models the acoustic token in an autoregressive way conditioned on the semantic token. VALL-E [6], the first TTS framework to leverage large, diverse, and multi-speaker speech data, demonstrates strong in-context learning capabilities similar to GPT-3, treating TTS as a language modeling task on audio codecs. MusicGen [12] generates music using a single-stage transformer LM alongside efficient token interleaving patterns. Similarly, UniAudio [7] scaled up to 165K hours of audio and 1B parameters, utilizing LLM techniques to generate tokens for various types of audio, including speech, sounds, music, and singing, given different input conditions.
While these works have shown success in developing audio language models, they all rely on the acoustic codecs such as Encodec [10] or Soundstream [8] for audio tokenization and de-tokenization. However, these acoustic codecs were originally designed for audio compression rather than for audio language models. This misalignment means the design may not be optimal for audio language modeling.
@@ 1 Introduction
To address this issue, approaches like SpeechTokenizer [13] have attempted to disentangle speech into separate tokens for content and timbre and perform distillation-based semantic and acoustic integration. However, this method may not integrate smoothly with all audio LLMs, especially those requiring uniform token treatment across different layers, such as utilizing flattened codec tokens [7, 12].
@@
2 Related Works
2.1 Audio Language Model
The success of Large Language Models (LLMs) has sparked a significant trend in leveraging language foundation models for audio generation tasks [14, 15, 16, 17, 7, 18]. Audio, much like language, consists of variable-length sequences, making it well-suited for modeling with language foundation models. One pioneering method, AudioLM [5], employs a multi-stage strategy to harness the predictive capabilities of foundation models for generating tokens unconditionally. This approach involves predicting semantic tokens from various conditions (e.g., phonemes, text descriptions, MIDI) in the initial stage, followed by transforming them into acoustic tokens through coarse-to-fine modeling, ultimately generating the waveform. Representative systems such as SPEAR-TTS [19] for speech synthesis and MusicLM [4] for music generation have also been proposed. However, the two-stage process can lead to complexity in training and suboptimal performance due to the separate development of semantic and acoustic tokens, leading to error accumulation.
Conversely, recent advancements have shown that methods employing a single-stage language model outperform two-stage approaches. For example, VALL-E [6] utilizes an autoregressive (AR) model to predict the first token and a non-autoregressive (NAR) model to estimate the residual tokens, demonstrating superior performance compared to AudioLM. Similarly, MusicGen [12] employs a single-stage transformer language model and incorporates a delay pattern strategy for efficient token interleaving, achieving better results than MusicLM. Other notable works include CLAM-TTS [20], VoiceCraft [21], and UniAudio [7].
@@ 2.1 Audio Language Model
2.2 Audio Codec
Recent advancements have seen a surge in deep learning methodologies employing vector quantization [22] to reconstruct continuous signals into discrete representations for AR generation. Notably, audio codecs based on the VQ-GAN framework [23] have gained prominence. For example, SoundStream [8] introduces a versatile codec adaptable to various audio types, integrating Residual Vector Quantization (RVQ) and Generative Adversarial Network (GAN) to refine quantization and reconstruction. Similarly, Encodec [10] enhances compression through a multi-scale discriminator and a loss-balancing strategy alongside a language model. HiFi-Codec [11] employs Group-Residual Vector Quantization (GRVQ) to minimize the need for extensive codebooks while maintaining high reconstruction fidelity. DAC [9] addresses codebook collapse, where some codes remain unused, by applying improved codebook learning to achieve higher compression rates.
These codecs primarily focus on acoustic reconstruction and higher compression rates, often overlooking their potential as tokenizers for audio LLMs. Some attempts have been made to develop more suitable tokenizers for audio LLMs. For example, SpeechTokenizer [13] utilizes HuBERT to separate speech into distinct VQ components for content and timbre/acoustic details. This separation improves the modeling of content in the AR stage of VALL-E, while the NAR stage enriches the acoustic details. However, a distillation framework is exploited, this makes SpeechTokenizer may not be compatible with all LLM architectures, especially those that require uniform treatment of tokens, such as methods using flattened codec tokens [7, 12]. Another attempt is presented by SemantiCodec [24], which employs a pre-trained AudioMAE [25] to generate distinct semantic and acoustic tokens from mel-spectrograms. However, this method inherits the issues of SpeechTokenizer and introduces additional complexity in token modeling. Moreover, since the audioMAE is performed on 2D time-frequency mel-spectrograms, LLMs must effectively handle dual scales (time and frequency), which may require significant modifications to existing LLM structures.
@@ 3 Methods
3.1 Acoustic Audio codec
As illustrated in Figure 1, our model builds upon the framework established by existing acoustic codecs such as Encodec [10] and DAC[9]. An acoustic audio codec is composed of three main components: an acoustic encoder, a quantizer, and an acoustic decoder. The input of the codec is the raw waveform $\textbf{X}\in\mathbb{R}^{n}$, where $n$ represents the number of waveform samples. This waveform is fed into the acoustic encoder, which consists of several convolutional layers and employs temporal downscaling to extract frame-level latent acoustic features $\textbf{A}\in\mathbb{R}^{H_{a}\times T}$, where $H_{a}$ denotes the hidden size of the acoustic features and $T$ is the number of frames. These continuous features are then transformed into a series of discrete tokens $\textbf{Q}\in\mathbb{R}^{M\times T}$ using a Residual Vector Quantizer (RVQ) with $M$ quantizer layers. During training, a specific codebook for the quantizer is learned, enabling the conversion of discrete tokens back to continuous features $\textbf{A}_{q}\in\mathbb{R}^{H_{a}\times T}$. The acoustic decoder then reconstructs the waveform $\hat{\textbf{X}}$ from $\textbf{A}_{q}$ using several convolutional layers and temporal upsampling. The training process is supervised using various losses, including mel loss, STFT loss, and GAN loss, to ensure high-quality acoustic reconstruction.
3.2 Analysing Semantic Shortcoming
@@
3.3 Designing Auxiliary Semantic Module
Our approach employs a straightforward method that enhances audio codecs by directly concatenating semantic and acoustic features. Initially, we extract the semantic feature vector $\textbf{S}^{*}\in\mathbb{R}^{H_{s}\times T}$ from the audio waveform x. This extraction utilizes a self-supervised, pre-trained model such as HuBERT [26] or wav2vec 2.0 [27]. The extracted features are then processed through multiple convolutional layers within a semantic encoder to yield the refined semantic feature vector S. Concurrently, the acoustic branch produces the feature A. These outputs, S and A, are subsequently concatenated using a linear projection $\phi$, formulated as:
@@ 4.1.1 Baselines
4.1.2 Training Details of X-Codec
@@
4.1.3 Training Details of VALL-E
@@ 4.1.4 Evaluation Metrics
  • WER (Word Error Rate): We utilize an Automatic Speech Recognition (ASR) model to transcribe the generated audio [6]. The discrepancies between these transcriptions and the original texts are quantified using WER, providing a critical measure of audio intelligibility.
  • Sim-O (Similarity Objective): This metric assesses the objective similarity between synthesized speech and the original reference speech. Sim-O uses feature embeddings extracted from a pre-trained speaker verification model to measure this similarity [26, 20] [^7] https://github.com/microsoft/UniSpeech/tree/main/downstreams/speaker_verification, reflecting the codec’s ability to preserve speaker characteristics.
  • UTMOS: We evaluate the audio quality using UTMOS, a Speech MOS (Mean Opinion Score) predictor [28] [^8] https://github.com/tarepan/SpeechMOS that automatically measures the naturalness of speech. This metric provides insights into the overall auditory quality of the synthesized speech.
@@
Encodec [10] C+DNS+AS+FSD+J 47.17 0.09 1.24 6.37 0.33 3.02
DAC [9] C+DNS+V+AS+J+M 85.55 0.03 1.24 6.81 0.34 3.31
Speechtokenizer [13] LibriSpeech 7.53 0.10 1.26 5.24 0.36 3.84
@@
4.1.5 Zero-shot TTS Results
We use librispeech-test-clean [29]for zero-shot TTS evaluation following VALL-E-continual-setting [6]. The results in Table 1 demonstrate the following key findings:
@@ 4.1.5 Zero-shot TTS Results
hubert-ls-960 [26] - 3.3 4.1
Encodec [10] 1 21.5 28.3
Encodec [10] 8 17.5 27.0
DAC [9] 1 26.3 32.7
DAC [9] 12 21.7 33.2
Speechtoknizer [13] 1 3.5 4.3
Speechtoknizer[13] 8 3.6 4.5
@@
4.1.6 Analysing the Effect of Codec
To further analyse the above results caused by different audio codecs, we evaluate phonetic discriminability using the ABX error rate [30]. This metric assesses how well different codecs can distinguish between similar phonetic sounds within and across various contexts. We specifically examine the continuous representations for VQ as indicated by the results in the following table 2. We compare the performance of various models in terms of within and across phonetic discriminability:
@@ 4.1.6 Analysing the Effect of Codec
4.2 Music and Sound Generation
To the best of our knowledge, this is the first exploration into the potential benefits of incorporating semantic information into audio codecs for enhancing music and general sound generation through audio LLMs. Conventional methods for general audio representation learning, aiming at capturing the semantic discriminability of audios, are generally based on 2D mel-spectrogram, such as AudioMAE [25] and Beats [31]. These methods are in stark contrast to traditional codecs that process audio sequentially, frame-by-frame. This difference poses challenges for direct integration into existing audio generation frameworks.
@@ 4.2 Music and Sound Generation
4.2.1 Training Details of Self-Supervised General Audio Representation
HuBERT-General-Audio is trained using 8 NVIDIA H800 GPUs on 2.6 million tokens across 325,000 iterations. For training stability, we adopt an inverse square root learning schedule, a modification from the polynomial decay schedule originally utilized in [26]. The learning rate is set at 0.0003 with warmup steps of 32,000. Unlike the original HuBERT, which utilizes MFCCs as the training target unit designed specifically for speech, our model leverages the first VQ layer of Encodec as the training target for acoustic unit discovery in the general audio. This choice eliminates the need for the K-means discretization step, saving significant time and computational resources.
@@ 4.2.1 Training Details of Self-Supervised General Audio Representation
4.2.2 Music Continuation
Training Details: Acquiring high-quality text-music pair data is challenging; therefore, we gathered approximately 100,000 hours of music-only data, including about one million songs for the music continuation task. We deployed nanoGPT  [^9] https://github.com/karpathy/nanoGPT to implement a GPT-2-medium (approximately 300M parameters) [32] as our generative model. This model utilizes the first VQ from our codec to construct the training sequences, with additional experiments involving multiple VQs detailed in the appendix. We set the block size of sequence modelling to 4096, corresponding to roughly 82 seconds of audio, and adjust the vocabulary size from 50,257 to 1024, matching our codec’s codebook size. Other training hyperparameters are consistent with previous GPT-2-medium configurations. We train 300,000 steps on 8 NVIDIA H800 GPUs. The batch size is set to 20, with a learning rate of 3e-4 and a warmup phase of 2000 steps.
Experiments: For music continuation, we randomly crop 600 samples with each 40 seconds in duration from the MUSDB18 dataset [33]. The initial 10 seconds of each sample are used as prompts for the audio LLM, while the subsequent 30 seconds are generated by the model. These generated segments are then compared against the corresponding ground truth (GT) segments. To ensure that the assessment is independent of the codec’s reconstruction fidelity, both the generated and GT audio are reconstructed using the first VQ layer of the codec, ensuring performance differences attributed solely to the generative models themselves.
The evaluation metrics of the generated music include: Frechet Distance (FD) computed using features from Pretrained Audio Neural Networks (PANNs) [34], Frechet Audio Distance (FAD), and FD-MERT Layer 9 [35]. The results, as summarized in Table 7, reveal that the X-Codec significantly outperforms the baseline acoustic codec across all metrics. This superior performance indicates the X-Codec has a better understanding and enabling more effective reproduction of complex musical structures.
4.2.3 Text-to-Sound